OpenStudy (anonymous):

What is the sum of the geometric sequence -4, 28, -196, if there are 6 terms? My answer: -9,604

OpenStudy (anonymous):

what formula did you use?

hartnn (hartnn):

wht value of r you got ?

OpenStudy (anonymous):

an = 1 r n-1

hartnn (hartnn):

thats not correct formula for SUM

OpenStudy (anonymous):

\[an=(-4)(7)^{n-1}\] this is what i got

OpenStudy (anonymous):

OH do i use the arthemtic

OpenStudy (anonymous):

arithmetic

hartnn (hartnn):

no...use geometric only, but sum formula

OpenStudy (anonymous):

i have the formula now

OpenStudy (anonymous):

yes thanks

hartnn (hartnn):

\(\huge S_n = a_1 \dfrac{r^n-1}{r-1}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a1 = 1st term

hartnn (hartnn):

yes, so a1=... ? r=... ?

hartnn (hartnn):

you want sum of 6 tems, so n=6

OpenStudy (anonymous):

-4 = a1

hartnn (hartnn):

yes

hartnn (hartnn):

r=.... ?

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

a2/a1?

OpenStudy (anonymous):

divide 2nd term by 1st term*

OpenStudy (anonymous):

28/-4

OpenStudy (anonymous):

= -7

hartnn (hartnn):

yes, r=-7 now just plug in values.

OpenStudy (anonymous):

ok ill tell you my answer

OpenStudy (anonymous):

-42.8571428571

hartnn (hartnn):

no....

OpenStudy (anonymous):

plugged this in: -7(6)-1/-7 - 1

OpenStudy (anonymous):

-43/-8?

hartnn (hartnn):

you missed a1 =-4 also, don't forget about brackets -4 [(-7)^6-1] / [-7-1]

OpenStudy (anonymous):

5.375

OpenStudy (anonymous):

oh my error

hartnn (hartnn):

what is (-7)^6=... ?

OpenStudy (anonymous):

117649

OpenStudy (anonymous):

my calc doesn't allow me to plug in what you told me

hartnn (hartnn):

yes, so -4 [117648]/ (-8) = 117648/2 =... ?

OpenStudy (anonymous):

514706

OpenStudy (anonymous):

58824

OpenStudy (anonymous):

thank you:)))

hartnn (hartnn):

yes, i too get 58824

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

love u

hartnn (hartnn):

love you too <3